### v1

# Unconditional verification of quantum computation with classical light

Abstract
Verification of quantum computation is a task to efficiently check whether an output given from a quantum computer is correct. Existing verification protocols conducted between a quantum computer to be verified and a verifier necessitate quantum communication to unconditionally detect any malicious behavior of the quantum computer solving any promise problem in ${\sf BQP}$. In this paper, we remove the necessity of the communication of qubits by proposing a “physically-classical” verification protocol in which the verifier just sends coherent light to the quantum computer.
I Introduction
Quantum computers are expected to outperform classical computers in several applications, from cryptanalysis BHT97 ; S97 ; HSTX20 to physics K95 ; NPdB21 ; CDBBT24 and chemistry KJLMA08 ; GL22 ; CFGHLMW23 . On the flip side of these advantages, they are susceptible to noises. Therefore, to get benefits from quantum computers, it would be necessary to devise an efficient protocol for checking whether a quantum computer outputs a correct answer. This task is called verification of quantum computation GKK19 ; EHWRMPCK20 ; KR21 . Although someone may think that verification protocols become useless if a sufficient number of qubits for quantum error correction G20 will be realized, this is not the case because it will be still necessary to check whether an adopted quantum error correction scheme faithfully works. Multiple small-scale experiments BFKW13 ; ZZCPXYYYHXCLG20 ; JWQCCLXSZM20 ; FMMD21 have already been demonstrated toward the realization of verifiable quantum information processing.
Verification protocols are evaluated in terms of five properties: (i) whether the soundness is information theoretical or computational, (ii) what type of communication is required for a verifier, (iii) how many number of non-communicating provers, which are quantum computers to be verified, are necessary, (iv) the presence or absence of a trusted third party who surely follows procedures of the protocols, and (v) to which problems the protocols can be applied. Here, the information-theoretical and computational soundnesses mean that any computationally-unbounded and quantum-polynomial-time malicious provers who output incorrect answers can be rejected, respectively, and hence the former is a stronger property than the latter. As for (iii), it would, in general, be hard to guarantee that multiple provers do not communicate each other. Therefore, the ultimate goal is to devise a protocol such that (i) the soundness is information theoretical, i.e., even if the quantum computer outputs an incorrect answer by taking superpolynomial time, it can be properly detected, (ii) the classical communication is sufficient, (iii) & (iv) a single prover is sufficient, and (v) the protocol can be applied to any problem in ${\sf BQP}$, which is a set of promise problems (i.e., problems that can be answered by YES or NO) solvable in quantum polynomial time. However, it is hard to construct such outstanding protocol with the currently known theoretical techniques. Although its impossibility was not shown, which immediately implies ${\sf P}\neq{\sf PSPACE}$, some existing results FHM18 ; MT20 reveal the difficulty of its construction.
This situation leaded to several verification protocols M12 ; RUV13 ; M14 ; GKW15 ; HM15 ; MNS16 ; M16 ; HKSE17 ; KW17 ; FK17 ; FHM18 ; M18 ; DOF18 ; TM18 ; B18 ; FKD18 ; LMNT18 ; MTN18 ; TMMMF19 ; CGJV19 ; G19 ; GV19 ; FKD19 ; LDTAF19 ; ZH19 ; M20 ; ACGH20 ; CCY20 ; LMKO21 ; CLLW22 ; KKLMO22 ; M22 ; Z22 ; GMP23 ; LZH23 each of which has own advantage and disadvantage. They fall into five types of approaches FN1 as summarized in Fig. 1. In terms of practicality, we stick to achieving the information-theoretical soundness and keeping the number of participants other than the verifier one. This is because it would be unclear how to guarantee that the prover’s computation is completed in polynomial time, that the multiple provers do not communicate each other, and that a third party does not cooperate with the prover. In this sense, from Fig. 1, quantum communication is necessary to verify any problem in ${\sf BQP}$ with existing practical verification protocols.
Figure 1: Summary of existing verification protocols. The third column indicates what type of communication is required for the verifier. In the fifth column, the cross and circle marks mean that a trusted third party is not necessary and necessary, respectively. In the sixth column, BQP means that the protocols can be applied to any problem in ${\sf BQP}$. The disadvantages of the verification protocols are indicated by bold red words. Our contribution is emphasized by the bold blue word.
In this paper, we remove the necessity of the communication of qubits by proposing a “physically-classical” verification protocol in which the verifier just sends coherent light to the quantum computer to be verified. The transmission of coherent light is the same as classical communication in the sense that classical bits are sent by using light in the real world, while they are absolutely different from the viewpoint of information theory. To obtain our result, we first modify the verification protocol in Ref. M20 such that a trusted third party is combined with the verifier, and hence communication of qubits becomes necessary for the verifier. Then, to remove the communication of qubits, we combine it with a technique of using coherent light as a substitute for qubits. So far, such technique has been developed in several quantum information processing tasks such as quantum key distribution (QKD) B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 , blind quantum computation DKL12 , and the demonstration of quantum advantage CKDK21 . All techniques except for that in Ref. DKL12 were developed for the transmission of classical bits. For example, random single-qubit states chosen from $\{|0\rangle,|1\rangle,|\pm\rangle\}$ with |±$|\pm\rangle\equiv(|0\rangle+|1\rangle)/\sqrt{2}$ are transmitted to share a secret bit string between two parties in the original protocol BB84 of QKD. This quantum communication is replaced with the communication of coherent light in Refs. B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 . However, the transmission of classical bits is insufficient for existing verification protocols. This is why we use the technique in Ref. DKL12 . Moreover, since its proof-of-principle experiment was already demonstrated over a distance of 100 km fiber JWHXSZZLYWLXZP19 , its adoption should be preferable from a practical point of view.
The remaining issue to be resolved is that the technique in Ref. DKL12 seems to be incompatible with existing verification protocols. The technique replaces the transmission of single-qubit states in a single plane of the Bloch sphere with that of coherent light, and it works when the single-qubit states are chosen uniformly at random. Although the single-qubit states used in the protocol of Ref. M20 are in the $x$-$z$ plane of the Bloch sphere, these qubits have the same random basis, i.e., the basis is not random among them. To fill in this gap, we further modify the protocol in Ref. M20 (for detail, see Protocol 1).
II Verification protocol with quantum communication
As the first step to construct our physically-classical verification protocol, we propose a verification protocol with communication of qubits by modifying the protocol in Ref. M20 . The purpose of our protocol is to verify quantum computation solving any problem in ${\sf BQP}$:
Definition 1 (BV97 )
A promise problem $L=(L_{\rm yes},L_{\rm no})$ is in ${\sf BQP}$ if and only if there exists a uniform family $\{U_{x}\}_{x}$ of polynomial-size quantum circuits such that when $x\in L_{\rm yes}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\geq 2/3$, and when $x\in L_{\rm no}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\leq 1/3$. Here, $I\equiv|0\rangle\langle 0|+|1\rangle\langle 1|$ is the two-dimensional identity operator, $n$ is a polynomial in $|x|$, and $|x|$ is the length of the instance $x$.
Simply speaking, ${\sf BQP}$ is a set of problems that can be efficiently solved with a universal quantum computer.
The protocol in Ref. M20 is based on the local Hamiltonian problem KSV02 . Let $L$ be a promise problem in ${\sf BQP}$. For any instance $x\in L$, we define the $N$-qubit Hamiltonian FN2
$H_{x}$$\equiv$ $\sum_{1\leq i<j\leq N}\cfrac{p_{ij}^{(x)}}{2}\times$ (1)
$\left(\cfrac{I^{\otimes N}+c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}+\cfrac{I^{\otimes N}+c_{ij}^{(x)}Y_{i}\otimes Y_{j}}{2}\right),$
where for all $i$, $j$, and $x$, $p_{ij}^{(x)}\geq 0$, $\sum_{i<j}p_{ij}^{(x)}=1$, $c_{ij}^{(x)}\in\{1,-1\}$, and $X_{i}$ and $Y_{i}$ represent the Pauli-$X$ and $Y$ operators applied to the $i$-th qubit, respectively. Note that $N$ is a polynomial in $|x|$. Let us assume that the prover declares $x\in L_{\rm yes}$, i.e., the quantum computer outputs that the correct answer is YES. We also define certain non-negative values $a$ and $b$ such that $1\geq b-a\geq 1/f(|x|)$ for a polynomial function $f$. From the ${\sf BQP}$-hardness (more precisely, ${\sf QMA}$-completeness) of the $2$-local Hamiltonian problem KKR06 ; BL08 ; CM16 , the verifier can efficiently find $\{p_{ij}^{(x)},c_{ij}^{(x)}\}_{1\leq i<j\leq N}$ such that (i) when the prover is honest (i.e., the correct answer is indeed YES), there exists an efficiently preparable quantum state $|\eta\rangle$ whose energy $\langle\eta|H_{x}|\eta\rangle$ is at most $a$, and (ii) when the prover is malicious (i.e., the correct answer is NO), the ground-state energy is at least $b$. Since ${\sf BQP}$ is closed under complement, even when the prover declares $x\in L_{\rm no}$, the verifier can efficiently find $\{p_{ij}^{(x)},c_{ij}^{(x)}\}_{1\leq i<j\leq N}$ having the same property. The verifier in the protocol of Ref. M20 decides whether the prover is honest or malicious by measuring the energy of $H_{x}$ with the aid of a trusted third party.
With the above idea in mind, we modify the protocol in Ref. M20 as follows:
[Protocol 1]
  1. 1.
    The verifier chooses two tuples $(h_{1},\ldots,h_{N})\in\{0,1\}^{N}$ and $(s_{1},\ldots,s_{N})\in\{0,1\}^{N}$ uniformly at random. Then the verifier sends
    $|\psi_{V}\rangle\equiv\bigotimes_{i=1}^{N}\left(S^{h_{i}}H|s_{i}\rangle\right)$ (2)
    to the prover, where $S\equiv|0\rangle\langle 0|+i|1\rangle\langle 1|$ is the $S$ gate, and $H\equiv|+\rangle\langle 0|+|-\rangle\langle 1|$ is the Hadamard gate.
  2. 2.
    The prover performs a POVM measurement $\{\Pi_{wz}\}_{w,z\in\{0,1\}^{N}}$ on the received state $|\psi_{V}\rangle$ and sends the measurement outcomes $w$ and $z$ to the verifier. If the prover is honest, $\{\Pi_{wz}\}_{w,z}$ corresponds to the $N$ Bell measurements on each qubits of $|\psi_{V}\rangle$ and $|\eta\rangle$. Therefore, the prover’s operation is essentially equivalent to the quantum teleportation of $|\eta\rangle$. On the other hand, if the prover is malicious, $\{\Pi_{wz}\}_{w,z}$ can be an arbitrary measurement.
  3. 3.
    The verifier samples a set $(i,j)$ with probability $p_{ij}^{(x)}$. Since the cardinality of the set $\{p_{ij}^{(x)}\}_{i<j}$ is $N(N-1)/2$, this sampling can be performed in classical polynomial time in $N$. If $h_{i}=h_{j}$, then the verifier proceeds to the next step. Otherwise, the verifier accepts the prover.
  4. 4.
    Let $s^{\prime}_{k}\equiv s_{k}\oplus z_{k}\oplus h_{k}w_{k}$ for all $1\leq k\leq N$, where $z_{k}$ and $w_{k}$ are the $k$-th bits of $z$ and $w$, respectively. If $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}=-c_{ij}^{(x)}$, the verifier accepts the prover. Otherwise, the verifier rejects the prover.
There exist two differences between the original protocol in Ref. M20 and Protocol 1. First, a trusted third party is merged with the verifier in step 1. Second, the bases $\{h_{i}\}_{i=1}^{N}$ are randomly chosen for each qubit in Protocol 1, while the basis of all qubits is determined by a single random bit $h\in\{0,1\}$ in the original protocol. These differences are essential to devise our physically-classical verification protocol in the next section.
As shown in Appendix A, the acceptance probability $p_{\rm acc}$ of Protocol 1 is at least $21-a/$ 2 (at most $21-b/$ 2) when the prover is honest (malicious). Since the gap of $p_{\rm acc}$ in the two cases is
$\left(1-\cfrac{a}{2}\right)-\left(1-\cfrac{b}{2}\right)=\cfrac{b-a}{2}\geq\cfrac{1}{2f(|x|)},$ (3)
the verifier can distinguish between the honest and malicious provers by repeating Protocol 1 in parallel a polynomial number of times.
III Verification of quantum computation with coherent light
The purpose of this section is to replace the quantum communication in Protocol 1 with the transmission of coherent light. To this end, we use the remote blind qubit state preparation (RBSP) protocol in Ref. DKL12 . The adoption of the RBSP protocol is possible due to the modification in the previous section.
We replace step 1 in Protocol 1 with the following protocol:
[Protocol 2]
  1. 1.
    The verifier and prover conduct the following steps $N$ times. Note that when the prover is malicious, the prover can apply any completely-positive trace-preserving (CPTP) map in each step.
    1. (a)
      Let $m(\geq 8)$ be a natural number specified later. For the $j$-th repetition $(1\leq j\leq N)$, the verifier chooses a tuple $(\theta_{1}^{(j)},\ldots,\theta_{m}^{(j)})\in\{0,\pi/2,\pi,3\pi/2\}^{m}$ uniformly at random. Then the verifier sends the $m$ phase-randomized coherent states
      $\bigotimes_{k=1}^{m}\left(e^{-\alpha^{2}}\sum_{n=0}^{\infty}\cfrac{\alpha^{2n}}{n!}|n_{\theta_{k}^{(j)}}\rangle\langle n_{\theta_{k}^{(j)}}|\right)$ (4)
      to the prover, where $(8/m)^{1/4}\leq\alpha\leq 1$, and $|n_{\theta}\rangle\equiv[(|0\rangle+e^{i\theta}|1\rangle)/\sqrt{2}]^{\otimes n}$ is an $n$-photon state with the polarization angle $\theta$. Here, $|0\rangle$ and $|1\rangle$ are not photon number states, but the computational basis ones.
    2. (b)
      The prover performs a quantum nondemolition (QND) measurement of the photon number on each of $m$ coherent states and obtains the measurement outcomes $\{n_{k}^{(j)}\}_{k=1}^{m}$. If $n_{k}^{(j)}\geq 1$, the prover keeps $|1_{\theta_{k}^{(j)}}\rangle$ at hand and discards the other $(n_{k}^{(j)}-1)$ photons. Let $m_{0}^{(j)}$ be the number of $k$’s such that $n_{k}^{(j)}=0$. At the end of this step, the honest prover possesses exactly $(m-m_{0}^{(j)})$ photons FNM1 .
    3. (c)
      The prover sends $\{n_{k}^{(j)}\}_{k=1}^{m}$ to the verifier.
    4. (d)
      The verifier calculates $m_{0}^{(j)}$ from $\{n_{k}^{(j)}\}_{k=1}^{m}$. If
      $m_{0}^{(j)}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right),$ (5)
      the verifier and prover proceed to the next step. Otherwise, the verifier rejects the prover.
    5. (e)
      The prover performs the interlaced 1D cluster computation (I1DC) protocol on the $(m-m_{0}^{(j)})$ photons as follows: for $l^{\prime}=1$ to $m-m_{0}^{(j)}-1$
      1. i.
        Apply $CZ(H\otimes I)$ to the $l^{\prime}$-th and $(l^{\prime}+1)$-th photons, where $CZ\equiv I^{\otimes 2}-2|11\rangle\langle 11|$ is the controlled-$Z$ gate.
      2. ii.
        Measure the $l^{\prime}$-th photon in the Pauli-$X$ basis, and obtain the measurement outcome $o_{l^{\prime}}^{(j)}$.
    6. (f)
      The prover sends the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}$ to the verifier. The prover keeps the unmeasured $(m-m_{0}^{(j)})$-th photon at hand and will use it as the $j$-th qubit $|\psi_{V}^{(j)}\rangle$ in step 2 of Protocol 1.
Figure 2: Schematic of our verification protocol with an honest prover. Classical communications and operations are represented by bold arrows. Non-classical communication and operations are represented by thin arrows. Each purple rectangular enclosure represents the Bell measurement. The verifier first sends phase-randomized coherent states with randomized polarization angles $\{\theta_{k}^{(j)}\}_{1\leq k\leq m,1\leq j\leq N}\in\{0,\pi/2,\pi,3\pi/2\}^{mN}$ to the prover. Then the prover performs QND measurements on all the received coherent states and sends the measurement outcomes $\{n_{k}^{(j)}\}_{k,j}$ to the verifier. The prover also discards unnecessary photons and performs the I1DC protocol by using the remaining photons (see steps (b) and (e) in Protocol 2). As a result, the prover obtains an $N$-qubit state $\otimes_{j=1}^{N}|\psi_{V}^{(j)}\rangle$, which will be used as $|\psi_{V}\rangle$ in step 1 of Protocol 1. The prover performs the quantum teleportation of the low-energy state $|\eta\rangle$ by measuring $N$ pairs of each qubits of $|\psi_{V}\rangle$ and $|\eta\rangle$ in the Bell bases and sending $\{w,z\}_{w,z\in\{0,1\}^{N}}$ to the verifier. On the other hand, the verifier checks if Eq. (5) holds for all $1\leq j\leq N$. If it does not hold, the verifier rejects the prover. If it holds, the prover chooses the pair $(i,j)$ with probability $p_{ij}^{(x)}$ and then checks whether $h_{i}=h_{j}$ (see step 3 in Protocol 1). If it is not satisfied, the prover is automatically accepted. Otherwise, the verifier calculates $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}$ by using $w$ and $z$ sent from the prover (see step 4 in Protocol 1). When it is equal to $-c_{ij}^{(x)}$, the verifier accepts the prover. On the other hand, it is not the case, the prover is rejected.
By replacing step 1 in Protocol 1 with Protocol 2, we can devise our physically-classical verification protocol (see Fig. 2). To evaluate its performance, we first derive its lower bound on the acceptance probability $p_{\rm acc}$ for the honest prover as a function of $\alpha$ and $m$. In step (b) of Protocol 2, the honest prover measures
$e^{-\alpha^{2}}\sum_{n=0}^{\infty}\cfrac{\alpha^{2n}}{n!}|n_{\theta_{k}^{(j)}}\rangle\langle n_{\theta_{k}^{(j)}}|$ (6)
in the Fock basis $\{|n\rangle\}_{n\in\mathbb{Z}_{\geq 0}}$. From Eq. (6), the probability of obtaining $n_{k}^{(j)}=0$ is $e^{-\alpha^{2}}$ for any $k$ and $j$, and hence the mean value of $m_{0}^{(j)}$ is $me^{-\alpha^{2}}$. The Hoeffding inequality H63 implies that the probability of the honest prover satisfying Eq. (5) is
${\rm Pr}\left[m_{0}^{(j)}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$ (7)
$\geq$ $1-{\rm Pr}\left[m_{0}^{(j)}-me^{-\alpha^{2}}\geq me^{-\alpha^{2}}\cfrac{\alpha^{2}}{2}\right]$
$\geq$ $1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (8)
Furthermore, when Eq. (5) is satisfied,
$m-m_{0}^{(j)}$$\geq$ $m-m\left(1-\cfrac{\alpha^{2}}{2}\right)\left(1+\cfrac{\alpha^{2}}{2}\right)$ (9)
$=$ $m\cfrac{\alpha^{4}}{4}\geq 2,$ (10)
and hence the prover can definitely perform the I1DC protocol in step (e) of Protocol 2.
For any $1\leq l\leq m-m_{0}^{(j)}$, let $\sigma_{l}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$ be the polarization angle of the $l$-th remaining photon at the end of step (b) in Protocol 2, i.e., the state of the $l$-th remaining photon is $(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}$ FNM3 . As shown in Ref. DKL12 , when the input states are $\{(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}\}_{l=1}^{m-m_{0}^{(j)}}$, the I1DC protocol outputs the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}\in\{0,1\}^{m-m_{0}^{(j)}-1}$ and the single-qubit state $(|0\rangle+e^{i\varphi^{(j)}}|1\rangle)/\sqrt{2}$, where
$\varphi^{(j)}\equiv\sum_{l=1}^{m-m_{0}^{(j)}-1}(-1)^{\sum_{l^{\prime}=l}^{m-m_{0}^{(j)}-1}o_{l^{\prime}}^{(j)}}\sigma_{l}^{(j)}+\sigma_{m-m_{0}^{(j)}}^{(j)}.$ (11)
Therefore, the verifier can calculate the value of $\varphi^{(j)}$ from the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}$ and the polarization angles $\{\sigma_{l}^{(j)}\}_{l=1}^{m-m_{0}^{(j)}}$ in classical polynomial time in $m$. The value of $\varphi^{(j)}$ is chosen from $\{0,\pi/2,\pi,3\pi/2\}$ with the same probability, $1/4$. This is because the verifier chooses the value of $\sigma_{m-m_{0}^{(j)}}^{(j)}$ uniformly at random in step (a) of Protocol 2. From the above argument, the output state of the I1DC protocol can be expressed as
$|\psi_{V}^{(j)}\rangle$ $=$ $S^{h_{j}}H|s_{j}\rangle$ (12)
by using two random bits $h_{j}$ and $s_{j}$. To be more specific, $(h_{j},s_{j})$ is $(0,0)$, $(0,1)$, $(1,0)$, or $(1,1)$ when $\varphi^{(j)}$ is 00, $\pi$, $\pi/2$, or $23\pi/$ 2, respectively.
In conclusion, if Eq. (5) is satisfied for all $1\leq j\leq N$, the verifier accepts the prover with the same probability as that of Protocol 1, i.e., with probability at least $21-a/$ 2. Thus, from Eq. (8), the lower bound on $p_{\rm acc}$ of our physically-classical verification protocol is
$\left[1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)\right]^{N}\left(1-\cfrac{a}{2}\right)$ (13)
$\geq$ $1-\cfrac{a}{2}-N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)$
when the prover is honest.
We next show the information-theoretical soundness of our physically-classical verification protocol. In other words, we give an upper bound on $p_{\rm acc}$ in the case of the malicious prover. To this end, we observe that the argument in Ref. DKL12 can be applied even in our situation. As an important point, how many photons are transmitted to the prover is randomly decided following the Poisson distribution, and hence the malicious prover cannot decide it even though any quantum operation is allowed for the malicious prover. We use this randomness to detect the malicious prover’s deviation.
Let $m_{0}^{(j)}$ and $m_{1}^{(j)}$ be the actual numbers of $k$’s such that $n_{k}^{(j)}=0$ and $1$, respectively. Although the malicious prover may not perform QND measurements in step (b) of Protocol 2, the actual photon number $n_{k}^{(j)}$ is properly defined because the phase-randomized coherent state in Eq. (6) is diagonalized in the Fock basis. Despite the malicious prover can perform any quantum operation, there exist only two cases where
  1. (i)
    for at least a single $j$, the number $\tilde{m}_{0}^{(j)}$ of vacuum states calculated from the measurement outcomes sent by the malicious prover in step (c) of Protocol 2 is at least $m_{0}^{(j)}+m_{1}^{(j)}$, and
  2. (ii)
    it is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$.
We derive an upper bound on the acceptance probability $p_{\rm acc}$ in each case one by one. By taking the maximum of these two upper bounds, we can derive an upper bound on $p_{\rm acc}$ of our physically-classical verification protocol.
We first consider the first case (i). Let $j^{\ast}$ such that $\tilde{m}_{0}^{(j^{\ast})}\geq m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$. From the Hoeffding inequality H63 , the probability of the prover not being rejected in step (d) of the $j^{\ast}$-th repetition of Protocol 2 is
${\rm Pr}\left[\tilde{m}_{0}^{(j^{\ast})}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$ (14)
$\leq$ ${\rm Pr}\left[m\cfrac{1+\alpha^{2}}{e^{\alpha^{2}}}-\left(m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}\right)\geq me^{-\alpha^{2}}\cfrac{\alpha^{2}}{2}\right]$
$\leq$ ${\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),$ (15)
where we have used $\tilde{m}_{0}^{(j^{\ast})}\geq m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$ and the fact that $m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$ converges to $me^{-\alpha^{2}}(1+\alpha^{2})$ to derive the first and second inequalities, respectively FNM2 . Since $\tilde{m}_{0}^{(j^{\ast})}\leq me^{-\alpha^{2}}(1+\alpha^{2}/2)$ is a necessary condition to be accepted, from Eq. (15),
$p_{\rm acc}\leq{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (16)
In the second case (ii), $\tilde{m}_{0}^{(j)}<m_{0}^{(j)}+m_{1}^{(j)}$ holds, and hence the input of the I1DC protocol must include at least a single state whose actual photon number is one or zero. Note that we can assume that the malicious prover is not rejected in step (d) of Protocol 2 because our purpose is to derive an upper bound on $p_{\rm acc}$.
We first consider the case where a single state whose actual photon number is one is included in the input of the I1DC protocol. Let it be the $l^{\ast}$-th input state whose polarization angle is $\sigma_{l^{\ast}}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$. In general, the malicious prover applies any CPTP map to the $(m-\tilde{m}_{0}^{(j)})$ input state that are used for the I1DC protocol in the case of the honest prover. We interpret it as that the $l^{\ast}$-th input state is converted to some quantum state by using the other $(m-\tilde{m}_{0}^{(j)}-1)$ input state as ancillary states. This interpretation implies that, immediately before step (e) in Protocol 2, the prover’s state is
$\cfrac{1}{4}\sum_{\sigma_{l^{\ast}}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}}\mathcal{E}\left(\cfrac{|0\rangle+e^{i\sigma_{l^{\ast}}^{(j)}}|1\rangle}{\sqrt{2}}\cfrac{\langle 0|+e^{-i\sigma_{l^{\ast}}^{(j)}}\langle 1|}{\sqrt{2}}\right)$ (17)
with some CPTP map $\mathcal{E}$. Since the verifier chooses each polarization angle independently in step (a) of Protocol 2, and the actual photon number of the $l^{\ast}$-th input state is one, $\mathcal{E}$ does not depend on $\sigma_{l^{\ast}}^{(j)}$ but can depend on the other input states’ polarization angles $\{\sigma_{l}^{(j)}\}_{1\leq l\leq m-\tilde{m}_{0}^{(j)},l\neq l^{\ast}}$. To mimic the honest prover, in steps (e) and (f) of Protocol 2, the malicious prover generates and sends the measurement outcomes $\vec{o}^{\ (j)}\equiv\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-\tilde{m}_{0}^{(j)}-1}\in\{0,1\}^{m-\tilde{m}_{0}^{(j)}-1}$ by applying some additional CPTP map to the quantum state in Eq. (17). For simplicity, we define $|+_{\theta}\rangle\equiv(|0\rangle+e^{i\theta}|1\rangle)/\sqrt{2}$ for any real number $\theta$. The malicious prover’s state is finally
$\cfrac{1}{4}\sum_{\sigma_{l^{\ast}}^{(j)}}\sum_{\vec{o}^{\ (j)}}p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})\mathcal{E}_{\vec{o}^{\ (j)}}\left(|+_{\sigma_{l^{\ast}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}}^{(j)}}|\right),$ (18)
where $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ is the probability of outputting $\vec{o}^{\ (j)}$, and $\mathcal{E}_{\vec{o}^{\ (j)}}$ is a quantum operation to be applied when the measurement outcomes are $\vec{o}^{\ (j)}$. The subscript of $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ just represents that the probability, in general, depends on $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$ and does not mean that the prover’s CPTP map is constructed by using the value of $\sigma_{l^{\ast}}^{(j)}$. In fact, there exits a CPTP map $\mathcal{F}_{1}^{(j)}$ such that
$\mathcal{F}_{1}^{(j)}\left(|+_{\sigma_{l^{\ast}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}}^{(j)}}|\right)$ (19)
$=$ $\sum_{\vec{o}^{\ (j)}}p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})\mathcal{E}_{\vec{o}^{\ (j)}}\left(|+_{\sigma_{l^{\ast}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}}^{(j)}}|\right)$
for any $\sigma_{l^{\ast}}^{(j)}$. This is due to the fact that the verifier keeps the value of $\sigma_{l^{\ast}}^{(j)}$ private, and the actual photon number of the $l^{\ast}$-th input state is one, i.e., the information of $\sigma_{l^{\ast}}^{(j)}$ is only contained in $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$. For any $\vec{o}^{\ (j)}$ and $\{\sigma_{l}^{(j)}\}_{l\neq l^{\ast}}$, there exist $\sigma\in\{0,\pi/2,\pi,3\pi/2\}$ and $c\in\{1,-1\}$ such that
$\varphi^{(j)}=\sigma+c\sigma_{l^{\ast}}^{(j)}.$ (20)
From Eq. (20), we can replace the variable $\sigma_{l^{\ast}}^{(j)}$ with $\varphi^{(j)}$ in Eq. (18) as follows:
$\cfrac{1}{4}\sum_{\varphi^{(j)}}\sum_{\vec{o}^{\ (j)}}p_{\varphi^{(j)}}(\vec{o}^{\ (j)})\mathcal{E}_{\vec{o}^{\ (j)}}\left(|+_{\varphi^{(j)}}\rangle\langle+_{\varphi^{(j)}}|\right)$ (21)
$=$ $\cfrac{1}{4}\sum_{\varphi^{(j)}}\sum_{\vec{o}^{\ (j)}}p_{\varphi^{(j)}}(\vec{o}^{\ (j)})\mathcal{E}_{\vec{o}^{\ (j)}}\left(|\psi_{V}^{(j)}\rangle\langle\psi_{V}^{(j)}|\right),$ (22)
where we have used the definition of $|\psi_{V}^{(j)}\rangle$ to obtain the equality. By applying Eq. (19) to Eq. (22), the malicious prover’s state at the end of Protocol 2 is
$\cfrac{1}{4}\sum_{\varphi^{(j)}}\mathcal{F}_{1}^{(j)}\left(|\psi_{V}^{(j)}\rangle\langle\psi_{V}^{(j)}|\right).$ (23)
This quantum state can also be prepared at the end of step 1 in Protocol 1.
We next consider the case where a vacuum state is included in the input of the I1DC protocol. Let it be originated from the $k^{\ast}$-th coherent state whose polarization angle is $\theta_{k^{\ast}}^{(j)}$. Since the verifier selects each polarization angle independently, the malicious prover’s state $\rho^{(j)}$ at the end of Protocol 2 does not depend on $\theta_{k^{\ast}}^{(j)}$ and hence $\varphi^{(j)}$. Remember that $\varphi^{(j)}$ is just a variable replacement of $\theta_{k^{\ast}}^{(j)}$. By using the CPTP map $\mathcal{F}_{2}^{(j)}$ that replaces any quantum state with the fixed state $\rho^{(j)}$, the prover’s final state is
$\mathcal{F}_{2}^{(j)}\left(|\psi_{V}^{(j)}\rangle\langle\psi_{V}^{(j)}|\right).$ (24)
This quantum state can also be prepared at the end of step 1 in Protocol 1.
By combining the above arguments, when the deviation of the prover is independent in each repetition, the malicious prover’s state after the $N$-th repetition in Protocol 2 is
$\cfrac{1}{4^{N}}\sum_{\{\varphi^{(j)}\}_{j=1}^{N}}\bigotimes_{j=1}^{N}\mathcal{F}^{(j)}\left(|\psi_{V}^{(j)}\rangle\langle\psi_{V}^{(j)}|\right)$ (25)
$=$ $\cfrac{1}{4^{N}}\sum_{\{\varphi^{(j)}\}_{j=1}^{N}}\left(\prod_{j=1}^{N}\mathcal{F}^{(j)}\right)\left(|\psi_{V}\rangle\langle\psi_{V}|\right),$ (26)
where $\mathcal{F}^{(j)}\in\{\mathcal{F}_{1}^{(j)},\mathcal{F}_{2}^{(j)}\}$ for all $1\leq j\leq N$. As shown in Appendix B, the similar argument holds even when the prover’s attack is collective, i.e., the prover simultaneously handles all photons in all repetitions to deceive the verifier. Therefore, $p_{\rm acc}\leq 1-b/2$. From this upper bound and Eq. (16), when the prover is malicious,
$p_{\rm acc}\leq{\rm max}\left\{{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),1-\cfrac{b}{2}\right\}.$ (27)
[…]
$1-\cfrac{a}{2}-N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)$ (28)
$-{\rm max}\left\{{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),1-\cfrac{b}{2}\right\}$
$\geq$ $1-\cfrac{a}{2}-\cfrac{1}{2Nf(|x|)}-\left(1-\cfrac{b}{2}\right)$
$\geq$ $\cfrac{N-1}{2Nf(|x|)},$ (29)
which is the inverse of a polynomial in $|x|$. Thus, our physically-classical verification protocol efficiently distinguishes the honest and malicious provers.
IV Conclusion & discussion
We have proposed an efficient verification protocol that removes the necessity of the communication of qubits. Since all apparatuses required for the verifier are a telecom-band laser with linear optical elements and a classical computer, our results would facilitate the realization of verifiable quantum computers.
To further improve the practicality of our protocol, we discuss the phase randomization implemented for the coherent state. In several papers (e.g., Ref. CZLM15 ), the difficulty of the continuous phase randomization was pointed out. By calculating the fidelity between the continuous-phase-randomized (see Eq. (4)) and discrete-phase-randomized coherent states, we evaluate how many classical bits would be required for the phase randomization. For simplicity, let
$\rho_{\theta_{k}^{(j)}}^{\infty}\equiv e^{-\alpha^{2}}\sum_{n=0}^{\infty}\cfrac{\alpha^{2n}}{n!}|n_{\theta_{k}^{(j)}}\rangle\langle n_{\theta_{k}^{(j)}}|$ (30)
and
$\rho_{\theta_{k}^{(j)}}^{R}\equiv\cfrac{1}{R}\sum_{j=0}^{R-1}|{e^{i2j\pi/R}\alpha}_{\theta_{k}^{(j)}}\rangle\langle{e^{i2j\pi/R}\alpha}_{\theta_{k}^{(j)}}|$ (31)
be the continuous-phase-randomized and discrete-phase-randomized coherent states, where $|\beta_{\theta}\rangle\equiv e^{-|\beta|^{2}/2}\sum_{n=0}^{\infty}(\beta^{n}/\sqrt{n!})|n_{\theta}\rangle$ is the phase-fixed coherent state with the polarization angle $\theta$ for any complex number $\beta$. We denote the fidelity between two quantum states $\rho$ and $\sigma$ as $F(\rho,\sigma)$. From Ref. JYWCGH24 , when $\alpha=1$ and $R\geq e^{2}+1$, the fidelity is
$F\left(\bigotimes_{j,k}\left(\rho_{\theta_{k}^{(j)}}^{\infty}\otimes|\theta_{k}^{(j)}\rangle\langle\theta_{k}^{(j)}|\right),\bigotimes_{j,k}\left(\rho_{\theta_{k}^{(j)}}^{R}\otimes|\theta_{k}^{(j)}\rangle\langle\theta_{k}^{(j)}|\right)\right)$ (32)
$=$ $F\left(\rho_{\theta_{k}^{(j)}}^{\infty},\rho_{\theta_{k}^{(j)}}^{R}\right)^{mN}$
$=$ $\left\{\sum_{j=0}^{R-1}\sqrt{\sum_{k=0}^{\infty}\left[\cfrac{e^{-1}}{(kR+j)!}\right]^{2}}\right\}^{2mN}$ (33)
$\geq$ $e^{-2mN}\left(\sum_{j=0}^{R-1}\cfrac{1}{j!}\right)^{2mN}$ (34)
$\geq$ $e^{-2mN}\left[e-\cfrac{e}{(R-1)!}\right]^{2mN}$ (35)
$\geq$ $1-2mN\left(\cfrac{e}{R-1}\right)^{R-1}\equiv F_{\rm min}.$ (36)
Let $p_{\rm acc}$ and $q_{\rm acc}$ be the acceptance probabilities of our physically-classical verification protocol with continuous-phase-randomized and discrete-phase-randomized coherent states, respectively. Eq. (29) implies that $|p_{\rm acc}-q_{\rm acc}|\leq(N-1)/[4Nf(|x|)]$ is sufficient to that our protocol correctly distinguishes the honest and malicious provers. From Ref. HQ12 that studies the fidelity and trace distance for any quantum states in an infinite-dimensional separable complex Hilbert space, $|p_{\rm acc}-q_{\rm acc}|\leq\sqrt{1-F_{\rm min}}$, and hence using $R\geq e^{2}+1$ results in $R=\lceil\log{[32mN^{3}f(|x|)^{2}/(N-1)^{2}]}+1\rceil$. Since $m=\lceil 2e^{2}\log{(2N^{2}f(|x|))}\rceil$, this calculation shows that a logarithmic of logarithmic number of classical bits is sufficient for the phase randomization.
The RBSP protocol DKL12 was improved or modified in Refs. ZL16 ; ZL18 ; NHS19 ; PYZMDZLW22 . As another direction to improve our protocol, it would be interesting to consider their applicability to the verification of quantum computation. Furthermore, as with the implementation security ZNC24 of QKD, it would be important to devise the verification protocols under several imperfections such as the channel loss and noises and the correlation between coherent states. An efficient way KLMO24 that removes trusted quantum state preparations and measurements from verification protocols may be useful for this purpose.
In this paper, we propose a physically-classical verification protocol by combining the protocols in Refs. M20 and DKL12 . On the other hand, there are compilers that make blind quantum computing protocols verifiable M18 ; KKLMO22 . Blind quantum computation is a secure protocol such that a user can delegate universal quantum computation to a remote quantum computer without disclosing the user’s input, quantum algorithm, and output. Although Ref. M18 seems to implicitly assume the perfect blindness (i.e., the perfect security), the blind quantum computing protocol with coherent states proposed in Ref. DKL12 does not satisfy the perfect blindness. This is why it would not be directly applied to Ref. DKL12 to obtain a physically-classical verification protocol. When we apply Ref. KKLMO22 to Ref. DKL12 , the resultant protocol should require a polynomial number of communication rounds between the verifier and prover, while our protocol is one round because $N$ repetitions in Protocol 2 can be done in parallel, and the prover can send $\{n_{k}^{(j)}\}_{k,j}$ and $\{w,z\}_{w,z}$ simultaneously. Our protocol successfully reveals an advantage of the transmission of coherent light in the sense that if the ultimate protocol in Fig. 1 with a constant number of rounds can be constructed, then ${\sf BQP}$ is contained in the third level of the polynomial hierarchy FHM18 . This set containment is considered to be unlikely from an oracle separation between ${\sf BQP}$ and ${\sf PH}$ RT22 .
It would be interesting to devise physically-classical verification protocols by modifying other existing verification protocols (e.g., Ref. B18 ) and then combining it with the RBSP protocol DKL12 .
ACKNOWLEDGMENTS
We thank Seiichiro Tani and Tomoyuki Morimae for helpful discussions. YT is partially supported by JST [Moonshot R&D – MILLENNIA Program] Grant Number JPMJMS2061. AM is partially supported by JST, ACT-X Grant No. JPMJAX210O, Japan.
Appendix A: Acceptance probabilities in Protocol 1
The calculation is based on the idea in Ref. MT20 . We first consider the case where the prover is honest. Let $h\equiv h_{1}\ldots h_{N}$, $s\equiv s_{1}\ldots s_{N}$, $\bar{\alpha}\equiv\alpha\oplus 1$ for any classical bit $\alpha\in\{0,1\}$, and $\delta_{\alpha\beta}$ be the Kronecker delta such that it is equal to one or zero when $\alpha=\beta$ or $\alpha\neq\beta$ for the two classical bits $\alpha\in\{0,1\}$ and $\beta\in\{0,1\}$, respectively. Since there exists the low-energy state $|\eta\rangle$, and $\{\Pi_{wz}\}_{w,z}$ corresponds to the $N$ parallel measurements in the Bell basis $\{|\phi_{\alpha\beta}\rangle\equiv(Z^{\beta}X^{\alpha}\otimes I)(|00\rangle+|11\rangle)/\sqrt{2}\}_{\alpha,\beta\in\{0,1\}}$, the acceptance probability $p_{\rm acc}$ is
$\cfrac{1}{2^{2N}}\sum_{h,s\in\{0,1\}^{N}}\sum_{w,z\in\{0,1\}^{N}}\left(\bigotimes_{k=1}^{N}\langle\phi_{w_{k}z_{k}}|\right)\left(|\eta\rangle\langle\eta|\otimes|\psi_{V}\rangle\langle\psi_{V}|\right)\left(\bigotimes_{k=1}^{N}|\phi_{w_{k}z_{k}}\rangle\right)\left\{\sum_{i<j}p_{ij}^{(x)}\left[\delta_{\bar{h_{i}}h_{j}}+\delta_{h_{i}h_{j}}\cfrac{1-c_{ij}^{(x)}(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}}{2}\right]\right\}$ (38)
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{2N}}\sum_{h,s,w,z\in\{0,1\}^{N}}\left(\bigotimes_{k=1}^{N}\langle\phi_{w_{k}z_{k}}|\right)\left(|\eta\rangle\langle\eta|\otimes|\psi_{V}\rangle\langle\psi_{V}|\right)\left(\bigotimes_{k=1}^{N}|\phi_{w_{k}z_{k}}\rangle\right)\left[\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\cfrac{1-c_{ij}^{(x)}(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}}{2}\right]$
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{3N}}\sum_{h,s,w,z\in\{0,1\}^{N}}\left(\bigotimes_{k=1}^{N}\langle s_{k}|H_{k}S_{k}^{h_{k}}X_{k}^{w_{k}}Z_{k}^{z_{k}}\right)|\eta\rangle\langle\eta|\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}}X_{k}^{w_{k}}{S_{k}^{\dagger}}^{h_{k}}H_{k}|s_{k}\rangle\right)\left[\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\cfrac{1-c_{ij}^{(x)}(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}}{2}\right]$ (42)
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{3N}}\sum_{h,s,w,z\in\{0,1\}^{N}}\left(\bigotimes_{k=1}^{N}\langle s_{k}|H_{k}S_{k}^{h_{k}}X_{k}^{w_{k}}Z_{k}^{z_{k}}\right)|\eta\rangle\langle\eta|\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}}X_{k}^{w_{k}}{S_{k}^{\dagger}}^{h_{k}}\right)\times$
$\left[\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\right]\left(\bigotimes_{k=1}^{N}H_{k}|s_{k}\rangle\right)$
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{3N}}\sum_{h,w,z\in\{0,1\}^{N}}\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\times$
${\rm Tr}\left[\left(\bigotimes_{k=1}^{N}S_{k}^{h_{k}}X_{k}^{w_{k}}Z_{k}^{z_{k}}\right)|\eta\rangle\langle\eta|\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}}X_{k}^{w_{k}}{S_{k}^{\dagger}}^{h_{k}}\right)\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\right]$
$=$ $\cfrac{1}{2}+\cfrac{1}{16}\sum_{i<j}\sum_{h_{i},w_{i},w_{j}\in\{0,1\}}p_{ij}^{(x)}\times$
${\rm Tr}\left[|\eta\rangle\langle\eta|\left(X_{i}^{w_{i}}{S_{i}^{\dagger}}^{h_{i}}\otimes X_{j}^{w_{j}}{S_{j}^{\dagger}}^{h_{i}}\right)\left(Z_{i}^{h_{i}w_{i}}\otimes Z_{j}^{h_{i}w_{j}}\right)\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}\left(Z_{i}^{h_{i}w_{i}}\otimes Z_{j}^{h_{i}w_{j}}\right)\left(S_{i}^{h_{i}}X_{i}^{w_{i}}\otimes S_{j}^{h_{i}}X_{j}^{w_{j}}\right)\right]$
$=$ $\cfrac{1}{2}+\cfrac{1}{2}{\rm Tr}\left[|\eta\rangle\langle\eta|\sum_{i<j}\cfrac{p_{ij}^{(x)}}{2}\left(\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}+\cfrac{I^{\otimes N}-c_{ij}^{(x)}Y_{i}\otimes Y_{j}}{2}\right)\right]$ (43)
$=$ $\cfrac{1+\langle\eta|\left(I^{\otimes N}-H_{x}\right)|\eta\rangle}{2}\geq 1-\cfrac{a}{2}.$ (44)
We next consider the case where the prover is malicious. The acceptance probability $p_{\rm acc}$ is
$\cfrac{1}{2^{2N}}\sum_{h,s\in\{0,1\}^{N}}\sum_{w,z\in\{0,1\}^{N}}\langle\psi_{V}|\Pi_{wz}|\psi_{V}\rangle\left\{\sum_{i<j}p_{ij}^{(x)}\left[\delta_{\bar{h_{i}}h_{j}}+\delta_{h_{i}h_{j}}\cfrac{1-c_{ij}^{(x)}(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}}{2}\right]\right\}$ (45)
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{2N}}\sum_{h,s,w,z\in\{0,1\}^{N}}\langle\psi_{V}|\Pi_{wz}|\psi_{V}\rangle\left\{\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\cfrac{1-c_{ij}^{(x)}(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}}{2}\right\}$ (49)
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{2N}}\sum_{h,s,w,z\in\{0,1\}^{N}}\left(\bigotimes_{k=1}^{N}\langle s_{k}|H_{k}{S_{k}^{\dagger}}^{h_{k}}\right)\Pi_{wz}\left(\bigotimes_{k=1}^{N}S_{k}^{h_{k}}\right)\times$
$\left[\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\right]\left(\bigotimes_{k=1}^{N}H_{k}|s_{k}\rangle\right)$
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{2N}}\sum_{h,w,z\in\{0,1\}^{N}}\times$
${\rm Tr}\left[\Pi_{wz}\left[\sum_{i<j}p_{ij}^{(x)}\delta_{h_{i}h_{j}}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\left(\bigotimes_{k=1}^{N}S_{k}^{h_{k}}\right)\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}\left(\bigotimes_{k=1}^{N}{S_{k}^{\dagger}}^{h_{k}}\right)\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}+h_{k}w_{k}}\right)\right]\right]$
$=$ $\cfrac{1}{2}+\cfrac{1}{2^{N+2}}\sum_{i<j}p_{ij}^{(x)}\sum_{h_{i}\in\{0,1\}}\sum_{w,z\in\{0,1\}^{N}}{\rm Tr}\left[\left(\bigotimes_{k=1}^{N}X_{k}^{w_{k}}Z_{k}^{z_{k}}\right)\Pi_{wz}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}}X_{k}^{w_{k}}\right)\times\right.$
$\left.\left[\left(X_{i}^{w_{i}}Z_{i}^{h_{i}w_{i}}\otimes X_{j}^{w_{j}}Z_{j}^{h_{i}w_{j}}\right)\left(S_{i}^{h_{i}}\otimes S_{j}^{h_{i}}\right)\cfrac{I^{\otimes N}-c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}\left({S_{i}^{\dagger}}^{h_{i}}\otimes{S_{j}^{\dagger}}^{h_{i}}\right)\left(Z_{i}^{h_{i}w_{i}}X_{i}^{w_{i}}\otimes Z_{j}^{h_{i}w_{j}}X_{j}^{w_{j}}\right)\right]\right]$
$=$ $\cfrac{1}{2}+\cfrac{1}{2}{\rm Tr}\left[\left[\cfrac{1}{2^{N}}\sum_{w,z\in\{0,1\}^{N}}\left(\bigotimes_{k=1}^{N}X_{k}^{w_{k}}Z_{k}^{z_{k}}\right)\Pi_{wz}\left(\bigotimes_{k=1}^{N}Z_{k}^{z_{k}}X_{k}^{w_{k}}\right)\right]\left(I^{\otimes N}-H_{x}\right)\right]\leq 1-\cfrac{b}{2},$ (50)
where we have used the observation that $[\sum_{w,z\in\{0,1\}^{N}}(\otimes_{k=1}^{N}X_{k}^{w_{k}}Z_{k}^{z_{k}})\Pi_{wz}(\otimes_{k=1}^{N}Z_{k}^{z_{k}}X_{k}^{w_{k}})]/2^{N}$ is a quantum state to obtain the last inequality.
Appendix B: Malicious prover’s final state in Protocol 2 for collective attacks
In this appendix, we derive the malicious prover’s quantum state after the $N$-th repetition in Protocol 2 under any collective attack. Since we consider the second case (ii) where the number $\tilde{m}_{0}^{(j)}$ of vacuum states is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$, the I1DC protocol in the $j$-th repetition has a state whose actual photon number is zero or one as an input. Let it be the $l^{\ast}_{j}$-th input state whose polarization angle is $\sigma_{l^{\ast}_{j}}^{(j)}$. For simplicity, we define $\vec{\sigma}\in\{0,\pi/2,\pi,3\pi/2\}^{(m-1)N}$ as the string of all the polarization angles except for $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. In general, the prover’s final state can be written as
$\cfrac{1}{4^{mN}}\sum_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}\sum_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)}}p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}\left(\bigotimes_{j=1}^{N}|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}_{j}}^{(j)}}|\right),$ (51)
where $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ is the probability of outputting $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$, and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ is a quantum operation to be applied when the measurement outcomes are $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$. The subscript $\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ just represents that the probability can depend on each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ if the $l^{\ast}_{j}$-th input qubit of the I1DC protocol in the $j$-th repetition is not a vacuum state, and it does not mean that the prover’s CPTP map is constructed by using the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. Another subscript $\vec{\sigma}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ means that they may depend on $\vec{\sigma}$ because the corresponding coherent states may include more than one photon. Since the verifier keeps the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$ private, and the actual photon number of each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is one or zero, as with Eq. (19), there exists a CPTP map $\mathcal{F}$ such that
$\mathcal{F}\left(\bigotimes_{j=1}^{N}|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}_{j}}^{(j)}}|\right)$ (52)
$=$ $\cfrac{1}{4^{(m-1)N}}\sum_{\vec{\sigma}}\sum_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)}}p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}\left(\bigotimes_{j=1}^{N}|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}_{j}}^{(j)}}|\right)$
for any $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$. Remember that when the actual photon number of $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is zero, the probability $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ does not depend on $\sigma_{l^{\ast}_{j}}^{(j)}$.
For any $\vec{\sigma}$ and $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$, there exist $\{\theta_{j}\}_{j=1}^{N}\in\{0,\pi/2,\pi,3\pi/2\}^{N}$ and $\{c_{j}\}_{j=1}^{N}\in\{1,-1\}^{N}$ such that
$\varphi^{(j)}=\theta_{j}+c_{j}\sigma_{l^{\ast}_{j}}^{(j)}$ (53)
for all $j$. It is worth mentioning that when the value of $\sigma_{l^{\ast}_{j}}^{(j)}$ is chosen from $\{0,\pi/2,\pi,3\pi/2\}$ uniformly at random, the value of $\varphi^{(j)}$ is also determined uniformly at random on the same range. Therefore, from Eqs. (53) and (52), the prover’s final state is
$\cfrac{1}{4^{mN}}\sum_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}\sum_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)}}p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}\left(\bigotimes_{j=1}^{N}|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}_{j}}^{(j)}}|\right)$ (55)
$=$ $\cfrac{1}{4^{mN}}\sum_{\varphi^{(1)},\varphi^{(2)},\ldots,\varphi^{(N)},\vec{\sigma}}\sum_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)}}p_{\varphi^{(1)},\varphi^{(2)},\ldots,\varphi^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}\left(\bigotimes_{j=1}^{N}|+_{\varphi^{(j)}}\rangle\langle+_{\varphi^{(j)}}|\right)$
$=$ $\cfrac{1}{4^{N}}\sum_{\varphi^{(1)},\varphi^{(2)},\ldots,\varphi^{(N)}}\mathcal{F}\left(\bigotimes_{j=1}^{N}|+_{\varphi^{(j)}}\rangle\langle+_{\varphi^{(j)}}|\right)$
$=$ $\cfrac{1}{4^{N}}\sum_{\varphi^{(1)},\varphi^{(2)},\ldots,\varphi^{(N)}}\mathcal{F}\left(|\psi_{V}\rangle\langle\psi_{V}|\right).$ (56)
Since this quantum state can be prepared at the end of step 1 in Protocol 1, the inequality $p_{\rm acc}\leq 1-b/2$ holds.

### v2

@@
 Existing verification protocols conducted between a quantum computer to be verified and a verifier necessitate quantum communication to unconditionally detect any malicious behavior of the quantum computer solving any promise problem in ${\sf BQP}$.
-In this paper, we remove the necessity of the communication of qubits by proposing a “physically-classical” verification protocol in which the verifier just sends coherent light to the quantum computer.
+In this paper, we remove the necessity of the communication of qubits by proposing a “physically classical” verification protocol in which the verifier just sends coherent light to the quantum computer.
 I Introduction
-Quantum computers are expected to outperform classical computers in several applications, from cryptanalysis BHT97 ; S97 ; HSTX20 to physics K95 ; NPdB21 ; CDBBT24 and chemistry KJLMA08 ; GL22 ; CFGHLMW23 .
-On the flip side of these advantages, they are susceptible to noises.
-Therefore, to get benefits from quantum computers, it would be necessary to devise an efficient protocol for checking whether a quantum computer outputs a correct answer.
-This task is called verification of quantum computation GKK19 ; EHWRMPCK20 ; KR21 .
-Although someone may think that verification protocols become useless if a sufficient number of qubits for quantum error correction G20 will be realized, this is not the case because it will be still necessary to check whether an adopted quantum error correction scheme faithfully works.
-Multiple small-scale experiments BFKW13 ; ZZCPXYYYHXCLG20 ; JWQCCLXSZM20 ; FMMD21 have already been demonstrated toward the realization of verifiable quantum information processing.
-Verification protocols are evaluated in terms of five properties: (i) whether the soundness is information theoretical or computational, (ii) what type of communication is required for a verifier, (iii) how many number of non-communicating provers, which are quantum computers to be verified, are necessary, (iv) the presence or absence of a trusted third party who surely follows procedures of the protocols, and (v) to which problems the protocols can be applied.
-Here, the information-theoretical and computational soundnesses mean that any computationally-unbounded and quantum-polynomial-time malicious provers who output incorrect answers can be rejected, respectively, and hence the former is a stronger property than the latter.
-As for (iii), it would, in general, be hard to guarantee that multiple provers do not communicate each other.
-Therefore, the ultimate goal is to devise a protocol such that (i) the soundness is information theoretical, i.e., even if the quantum computer outputs an incorrect answer by taking superpolynomial time, it can be properly detected, (ii) the classical communication is sufficient, (iii) & (iv) a single prover is sufficient, and (v) the protocol can be applied to any problem in ${\sf BQP}$, which is a set of promise problems (i.e., problems that can be answered by YES or NO) solvable in quantum polynomial time.
-However, it is hard to construct such outstanding protocol with the currently known theoretical techniques.
-Although its impossibility was not shown, which immediately implies ${\sf P}\neq{\sf PSPACE}$, some existing results FHM18 ; MT20 reveal the difficulty of its construction.
-This situation leaded to several verification protocols M12 ; RUV13 ; M14 ; GKW15 ; HM15 ; MNS16 ; M16 ; HKSE17 ; KW17 ; FK17 ; FHM18 ; M18 ; DOF18 ; TM18 ; B18 ; FKD18 ; LMNT18 ; MTN18 ; TMMMF19 ; CGJV19 ; G19 ; GV19 ; FKD19 ; LDTAF19 ; ZH19 ; M20 ; ACGH20 ; CCY20 ; LMKO21 ; CLLW22 ; KKLMO22 ; M22 ; Z22 ; GMP23 ; LZH23 each of which has own advantage and disadvantage.
-They fall into five types of approaches FN1 as summarized in Fig.
+Quantum computers are expected to outperform classical computers in a variety of applications, from cryptanalysis BHT97 ; S97 ; HSTX20 to physics K95 ; NPdB21 ; CDBBT24 and chemistry KJLMA08 ; GL22 ; CFGHLMW23 .
+The flip side of their advantages is their susceptibility to noises.
+Therefore, to obtain benefits from quantum computers, it is necessary to devise an efficient protocol for checking whether a quantum computer outputs a correct answer, a task called verification of quantum computation GKK19 ; EHWRMPCK20 ; KR21 .
+One may think that verification protocols would become useless if a sufficient number of qubits for quantum error correction G20 can be created; however, this is not the case, because it will still be necessary to check whether an implemented quantum error correction scheme faithfully works.
+Multiple small-scale experiments BFKW13 ; ZZCPXYYYHXCLG20 ; JWQCCLXSZM20 ; FMMD21 have already demonstrated progress toward the realization of verifiable quantum information processing.
+Verification protocols are evaluated in terms of five properties: (i) whether the soundness is information theoretic or computational; (ii) the type of communication required for a verifier; (iii) the number of necessary non-communicating provers, which are quantum computers to be verified; (iv) the presence or absence of a trusted third party who assuredly follows procedures in the protocols; and (v) the problems to which the protocols can be applied.
+Here, soundness can be information theoretic or computational.
+Information-theoretic soundness means that any malicious computationally unbounded prover who outputs incorrect answers can be rejected.
+On the other hand, computational soundness is a property for rejecting malicious quantum polynomial-time provers.
+Hence, the former property is stronger than the latter.
+As for (iii), it would, in general, be hard to guarantee that multiple provers do not communicate with each other.
+Therefore, the ultimate goal is to devise a protocol such that (i) the soundness is information theoretic, i.e., even if the quantum computer outputs an incorrect answer by taking superpolynomial time, it can be properly detected, (ii) the classical communication is sufficient, (iii) & (iv) a single prover is sufficient, and (v) the protocol is applicable to any problem in ${\sf BQP}$, which is a set of promise problems (i.e., problems that can be answered by YES or NO) solvable in quantum polynomial time.
+However, it is hard to construct such an outstanding protocol with currently known theoretical techniques.
+Although its impossibility has not been shown, which immediately implies ${\sf P}\neq{\sf PSPACE}$, some existing results FHM18 ; MT20 have revealed the difficulty of its construction.
+This situation has led to the development of numerous verification protocols M12 ; RUV13 ; M14 ; GKW15 ; HM15 ; MNS16 ; M16 ; HKSE17 ; KW17 ; FK17 ; FHM18 ; M18 ; DOF18 ; TM18 ; B18 ; FKD18 ; LMNT18 ; MTN18 ; TMMMF19 ; CGJV19 ; G19 ; GV19 ; FKD19 ; LDTAF19 ; ZH19 ; M20 ; ACGH20 ; CCY20 ; LMKO21 ; CLLW22 ; KKLMO22 ; M22 ; Z22 ; GMP23 ; LZH23 , each of which has its own advantages and disadvantages.
+These protocols fall into five types of approaches FN1 as summarized in Fig.
 1.
-In terms of practicality, we stick to achieving the information-theoretical soundness and keeping the number of participants other than the verifier one.
-This is because it would be unclear how to guarantee that the prover’s computation is completed in polynomial time, that the multiple provers do not communicate each other, and that a third party does not cooperate with the prover.
+In terms of practicality, we stick to achieving the information-theoretic soundness and keeping the total number of provers and a trusted third party one.
+This is because it would be unclear how to guarantee that the prover’s computation is completed in polynomial time, that the multiple provers do not communicate with each other, and that a third party does not cooperate with a prover.
 In this sense, from Fig.
@@
 Figure 1: Summary of existing verification protocols.
-The third column indicates what type of communication is required for the verifier.
-In the fifth column, the cross and circle marks mean that a trusted third party is not necessary and necessary, respectively.
+The third column indicates the type of communication required for the verifier.
+In the fifth column, the cross and circle marks mean that a trusted third party is unnecessary and necessary, respectively.
 In the sixth column, BQP means that the protocols can be applied to any problem in ${\sf BQP}$.
-The disadvantages of the verification protocols are indicated by bold red words.
-Our contribution is emphasized by the bold blue word.
-In this paper, we remove the necessity of the communication of qubits by proposing a “physically-classical” verification protocol in which the verifier just sends coherent light to the quantum computer to be verified.
+The disadvantages of the verification protocols are highlighted in bold red.
+Our contribution is emphasized in bold blue.
+In this paper, to remove the necessity of qubit communication, we propose a “physically classical” verification protocol in which the verifier just sends coherent light to the quantum computer to be verified.
 The transmission of coherent light is the same as classical communication in the sense that classical bits are sent by using light in the real world, while they are absolutely different from the viewpoint of information theory.
-To obtain our result, we first modify the verification protocol in Ref.
-M20 such that a trusted third party is combined with the verifier, and hence communication of qubits becomes necessary for the verifier.
-Then, to remove the communication of qubits, we combine it with a technique of using coherent light as a substitute for qubits.
-So far, such technique has been developed in several quantum information processing tasks such as quantum key distribution (QKD) B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 , blind quantum computation DKL12 , and the demonstration of quantum advantage CKDK21 .
-All techniques except for that in Ref.
+To obtain our protocol, we first modify the verification protocol in Ref.
+M20 such that a trusted third party is combined with the verifier, and hence qubit communication becomes necessary for the verifier.
+Then, to remove the qubit communication, we combine it with the technique of using coherent light as a substitute for qubits.
+This technique has been developed in several quantum information processing tasks such as quantum key distribution (QKD) B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 , blind quantum computation DKL12 , and the demonstration of quantum advantage CKDK21 .
+All the approaches except for that in Ref.
 DKL12 were developed for the transmission of classical bits.
-For example, random single-qubit states chosen from $\{|0\rangle,|1\rangle,|\pm\rangle\}$ with |±$|\pm\rangle\equiv(|0\rangle+|1\rangle)/\sqrt{2}$ are transmitted to share a secret bit string between two parties in the original protocol BB84 of QKD.
-This quantum communication is replaced with the communication of coherent light in Refs.
+For example, random single-qubit states chosen from $\{|0\rangle,|1\rangle,|\pm\rangle\}$ with |±$|\pm\rangle\equiv(|0\rangle+|1\rangle)/\sqrt{2}$ are transmitted to share a secret bit string between two parties in the original QKD protocol BB84 .
+This quantum communication is replaced with coherent light communication in Refs.
 B92 ; HIGM95 ; IWY03 ; LMC05 ; LP07 .
-However, the transmission of classical bits is insufficient for existing verification protocols.
-This is why we use the technique in Ref.
+However, the transmission of classical bits is insufficient for existing verification protocols, which is why we use the technique in Ref.
 DKL12 .
-Moreover, since its proof-of-principle experiment was already demonstrated over a distance of 100 km fiber JWHXSZZLYWLXZP19 , its adoption should be preferable from a practical point of view.
+Moreover, since its proof of principle experiment has already been demonstrated over a distance of 100 km fiber JWHXSZZLYWLXZP19 , its use is preferable from a practical point of view.
 The remaining issue to be resolved is that the technique in Ref.
@@
 To fill in this gap, we further modify the protocol in Ref.
-M20 (for detail, see Protocol 1).
+M20 (for details, see Protocol 1).
 II Verification protocol with quantum communication
-As the first step to construct our physically-classical verification protocol, we propose a verification protocol with communication of qubits by modifying the protocol in Ref.
+As the first step to constructing our physically classical verification protocol, we propose a verification protocol with communication of qubits by modifying the protocol in Ref.
 M20 .
@@
 Definition 1 (BV97 )
-A promise problem $L=(L_{\rm yes},L_{\rm no})$ is in ${\sf BQP}$ if and only if there exists a uniform family $\{U_{x}\}_{x}$ of polynomial-size quantum circuits such that when $x\in L_{\rm yes}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\geq 2/3$, and when $x\in L_{\rm no}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\leq 1/3$.
+A promise problem $L=(L_{\rm yes},L_{\rm no})$ is in ${\sf BQP}$ if and only if there exists a uniform family $\{U_{x}\}_{x}$ of polynomial-size quantum circuits such that when $x\in L_{\rm yes}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\geq 2/3$ and when $x\in L_{\rm no}$, $\langle 0^{n}|U_{x}^{\dagger}(|1\rangle\langle 1|\otimes I^{\otimes n-1})U_{x}|0^{n}\rangle\leq 1/3$.
 Here, $I\equiv|0\rangle\langle 0|+|1\rangle\langle 1|$ is the two-dimensional identity operator, $n$ is a polynomial in $|x|$, and $|x|$ is the length of the instance $x$.
@@
 $\left(\cfrac{I^{\otimes N}+c_{ij}^{(x)}X_{i}\otimes X_{j}}{2}+\cfrac{I^{\otimes N}+c_{ij}^{(x)}Y_{i}\otimes Y_{j}}{2}\right),$
-where for all $i$, $j$, and $x$, $p_{ij}^{(x)}\geq 0$, $\sum_{i<j}p_{ij}^{(x)}=1$, $c_{ij}^{(x)}\in\{1,-1\}$, and $X_{i}$ and $Y_{i}$ represent the Pauli-$X$ and $Y$ operators applied to the $i$-th qubit, respectively.
+where for all $i$, $j$, and $x$, $p_{ij}^{(x)}\geq 0$, $\sum_{i<j}p_{ij}^{(x)}=1$, $c_{ij}^{(x)}\in\{1,-1\}$, and $X_{i}$ and $Y_{i}$ represent the Pauli-$X$ and $Y$ operators applied to the $i$th qubit, respectively.
 Note that $N$ is a polynomial in $|x|$.
-Let us assume that the prover declares $x\in L_{\rm yes}$, i.e., the quantum computer outputs that the correct answer is YES.
+Let us assume that the prover declares $x\in L_{\rm yes}$, i.e., the quantum computer outputs YES as the correct answer.
 We also define certain non-negative values $a$ and $b$ such that $1\geq b-a\geq 1/f(|x|)$ for a polynomial function $f$.
@@
   4. 4.
-    Let $s^{\prime}_{k}\equiv s_{k}\oplus z_{k}\oplus h_{k}w_{k}$ for all $1\leq k\leq N$, where $z_{k}$ and $w_{k}$ are the $k$-th bits of $z$ and $w$, respectively.
+    Let $s^{\prime}_{k}\equiv s_{k}\oplus z_{k}\oplus h_{k}w_{k}$ for all $1\leq k\leq N$, where $z_{k}$ and $w_{k}$ are the $k$th bits of $z$ and $w$, respectively.
 If $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}=-c_{ij}^{(x)}$, the verifier accepts the prover.
@@
 Second, the bases $\{h_{i}\}_{i=1}^{N}$ are randomly chosen for each qubit in Protocol 1, while the basis of all qubits is determined by a single random bit $h\in\{0,1\}$ in the original protocol.
-These differences are essential to devise our physically-classical verification protocol in the next section.
+These differences are essential for devising our physically classical verification protocol, which is described in the next section.
 As shown in Appendix A, the acceptance probability $p_{\rm acc}$ of Protocol 1 is at least $21-a/$ 2 (at most $21-b/$ 2) when the prover is honest (malicious).
@@
 The adoption of the RBSP protocol is possible due to the modification in the previous section.
+Figure 2: Schematic of our verification protocol with an honest prover.
+Classical communication and operations are represented by bold arrows.
+Non-classical communication and operations are represented by thin arrows.
+Each purple rectangular enclosure represents the Bell measurement.
+The verifier first sends phase-randomized coherent states with randomized polarization angles $\{\theta_{k}^{(j)}\}_{1\leq k\leq m,1\leq j\leq N}\in\{0,\pi/2,\pi,3\pi/2\}^{mN}$ to the prover.
+Then the prover performs QND measurements on all the received coherent states and sends the measurement outcomes $\{n_{k}^{(j)}\}_{k,j}$ to the verifier.
+The prover also discards unnecessary photons and performs the I1DC protocol by using the remaining photons (see steps (b) and (e) in Protocol 2).
+As a result, the prover obtains an $N$-qubit state $\otimes_{j=1}^{N}|\psi_{V}^{(j)}\rangle$, which will be used as $|\psi_{V}\rangle$ in step 1 of Protocol 1.
+The prover performs the quantum teleportation of the low-energy state $|\eta\rangle$ by measuring $N$ pairs of qubits of $|\psi_{V}\rangle$ and $|\eta\rangle$ in the Bell bases and sending $\{w,z\}_{w,z\in\{0,1\}^{N}}$ to the verifier.
+On the other hand, the verifier checks whether Eq.
+(5) holds for all $1\leq j\leq N$.
+If it does not hold, the verifier rejects the prover.
+If it holds, the prover chooses the pair $(i,j)$ with probability $p_{ij}^{(x)}$ and then checks whether $h_{i}=h_{j}$ (see step 3 in Protocol 1).
+If it is not satisfied, the prover is automatically accepted.
+Otherwise, the verifier calculates $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}$ by using $w$ and $z$ sent from the prover (see step 4 in Protocol 1).
+When it is equal to $-c_{ij}^{(x)}$, the verifier accepts the prover.
+On the other hand, if it is not the case, the prover is rejected.
 We replace step 1 in Protocol 1 with the following protocol:
@@
   1. 1.
-    The verifier and prover conduct the following steps $N$ times. Note that when the prover is malicious, the prover can apply any completely-positive trace-preserving (CPTP) map in each step.
+    The verifier and prover conduct the following steps $N$ times. Note that when the prover is malicious, the prover can apply any completely positive trace-preserving (CPTP) map in each step.
     1. (a)
       Let $m(\geq 8)$ be a natural number specified later.
-For the $j$-th repetition $(1\leq j\leq N)$, the verifier chooses a tuple $(\theta_{1}^{(j)},\ldots,\theta_{m}^{(j)})\in\{0,\pi/2,\pi,3\pi/2\}^{m}$ uniformly at random.
+For the $j$th repetition $(1\leq j\leq N)$, the verifier chooses a tuple $(\theta_{1}^{(j)},\ldots,\theta_{m}^{(j)})\in\{0,\pi/2,\pi,3\pi/2\}^{m}$ uniformly at random.
 Then the verifier sends the $m$ phase-randomized coherent states
@@
     2. (b)
-      The prover performs a quantum nondemolition (QND) measurement of the photon number on each of $m$ coherent states and obtains the measurement outcomes $\{n_{k}^{(j)}\}_{k=1}^{m}$.
+      The prover performs a quantum nondemolition (QND) measurement of the photon number on each of the $m$ coherent states and obtains the measurement outcomes $\{n_{k}^{(j)}\}_{k=1}^{m}$.
 If $n_{k}^{(j)}\geq 1$, the prover keeps $|1_{\theta_{k}^{(j)}}\rangle$ at hand and discards the other $(n_{k}^{(j)}-1)$ photons.
@@
     5. (e)
-      The prover performs the interlaced 1D cluster computation (I1DC) protocol on the $(m-m_{0}^{(j)})$ photons as follows: for $l^{\prime}=1$ to $m-m_{0}^{(j)}-1$
+      The prover performs the interlaced 1D cluster computation (I1DC) protocol on the $(m-m_{0}^{(j)})$ photons as follows: for $l^{\prime}=1$ to $m-m_{0}^{(j)}-1$,
       1. i.
-        Apply $CZ(H\otimes I)$ to the $l^{\prime}$-th and $(l^{\prime}+1)$-th photons, where $CZ\equiv I^{\otimes 2}-2|11\rangle\langle 11|$ is the controlled-$Z$ gate.
+        Apply $CZ(H\otimes I)$ to the $l^{\prime}$th and $(l^{\prime}+1)$th photons, where $CZ\equiv I^{\otimes 2}-2|11\rangle\langle 11|$ is the controlled-$Z$ gate.
       2. ii.
-        Measure the $l^{\prime}$-th photon in the Pauli-$X$ basis, and obtain the measurement outcome $o_{l^{\prime}}^{(j)}$.
+        Measure the $l^{\prime}$th photon in the Pauli-$X$ basis, and obtain the measurement outcome $o_{l^{\prime}}^{(j)}$.
     6. (f)
       The prover sends the measurement outcomes $\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-m_{0}^{(j)}-1}$ to the verifier.
-The prover keeps the unmeasured $(m-m_{0}^{(j)})$-th photon at hand and will use it as the $j$-th qubit $|\psi_{V}^{(j)}\rangle$ in step 2 of Protocol 1.
-Figure 2: Schematic of our verification protocol with an honest prover.
-Classical communications and operations are represented by bold arrows.
-Non-classical communication and operations are represented by thin arrows.
-Each purple rectangular enclosure represents the Bell measurement.
-The verifier first sends phase-randomized coherent states with randomized polarization angles $\{\theta_{k}^{(j)}\}_{1\leq k\leq m,1\leq j\leq N}\in\{0,\pi/2,\pi,3\pi/2\}^{mN}$ to the prover.
-Then the prover performs QND measurements on all the received coherent states and sends the measurement outcomes $\{n_{k}^{(j)}\}_{k,j}$ to the verifier.
-The prover also discards unnecessary photons and performs the I1DC protocol by using the remaining photons (see steps (b) and (e) in Protocol 2).
-As a result, the prover obtains an $N$-qubit state $\otimes_{j=1}^{N}|\psi_{V}^{(j)}\rangle$, which will be used as $|\psi_{V}\rangle$ in step 1 of Protocol 1.
-The prover performs the quantum teleportation of the low-energy state $|\eta\rangle$ by measuring $N$ pairs of each qubits of $|\psi_{V}\rangle$ and $|\eta\rangle$ in the Bell bases and sending $\{w,z\}_{w,z\in\{0,1\}^{N}}$ to the verifier.
-On the other hand, the verifier checks if Eq.
-(5) holds for all $1\leq j\leq N$.
-If it does not hold, the verifier rejects the prover.
-If it holds, the prover chooses the pair $(i,j)$ with probability $p_{ij}^{(x)}$ and then checks whether $h_{i}=h_{j}$ (see step 3 in Protocol 1).
-If it is not satisfied, the prover is automatically accepted.
-Otherwise, the verifier calculates $(-1)^{s^{\prime}_{i}+s^{\prime}_{j}}$ by using $w$ and $z$ sent from the prover (see step 4 in Protocol 1).
-When it is equal to $-c_{ij}^{(x)}$, the verifier accepts the prover.
-On the other hand, it is not the case, the prover is rejected.
-By replacing step 1 in Protocol 1 with Protocol 2, we can devise our physically-classical verification protocol (see Fig.
+The prover keeps the unmeasured $(m-m_{0}^{(j)})$th photon at hand and will use it as the $j$th qubit $|\psi_{V}^{(j)}\rangle$ in step 2 of Protocol 1.
+By replacing step 1 in Protocol 1 with Protocol 2, we can devise our physically classical verification protocol (see Fig.
 2).
@@
 and hence the prover can definitely perform the I1DC protocol in step (e) of Protocol 2.
-For any $1\leq l\leq m-m_{0}^{(j)}$, let $\sigma_{l}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$ be the polarization angle of the $l$-th remaining photon at the end of step (b) in Protocol 2, i.e., the state of the $l$-th remaining photon is $(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}$ FNM3 .
+For any $1\leq l\leq m-m_{0}^{(j)}$, let $\sigma_{l}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$ be the polarization angle of the $l$th remaining photon at the end of step (b) in Protocol 2, i.e., the state of the $l$th remaining photon is $(|0\rangle+e^{i\sigma_{l}^{(j)}}|1\rangle)/\sqrt{2}$ FNM3 .
 As shown in Ref.
@@
 In conclusion, if Eq.
-(5) is satisfied for all $1\leq j\leq N$, the verifier accepts the prover with the same probability as that of Protocol 1, i.e., with probability at least $21-a/$ 2.
+(5) is satisfied for all $1\leq j\leq N$, the verifier accepts the prover with the same probability as that of Protocol 1, i.e., with probability of at least $21-a/$ 2.
 Thus, from Eq.
-(8), the lower bound on $p_{\rm acc}$ of our physically-classical verification protocol is
+(8), the lower bound on $p_{\rm acc}$ of our physically classical verification protocol is
 $\left[1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)\right]^{N}\left(1-\cfrac{a}{2}\right)$ (13)
@@
 when the prover is honest.
-We next show the information-theoretical soundness of our physically-classical verification protocol.
+We next show the information-theoretic soundness of our physically classical verification protocol.
 In other words, we give an upper bound on $p_{\rm acc}$ in the case of the malicious prover.
@@
 (6) is diagonalized in the Fock basis.
-Despite the malicious prover can perform any quantum operation, there exist only two cases where
+Despite the malicious prover’s ability to perform any quantum operation, there exist only two cases where
   1. (i)
-    for at least a single $j$, the number $\tilde{m}_{0}^{(j)}$ of vacuum states calculated from the measurement outcomes sent by the malicious prover in step (c) of Protocol 2 is at least $m_{0}^{(j)}+m_{1}^{(j)}$, and
+    for at least a single $j$, the actual number $m_{0}^{(j)}+m_{1}^{(j)}$ of vacuum and single-photon states is less than or equal to the threshold value $me^{-\alpha^{2}}(1+\alpha^{2}/2)$ in step (d) of Protocol 2, and
   2. (ii)
-    it is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$.
-We derive an upper bound on the acceptance probability $p_{\rm acc}$ in each case one by one.
-By taking the maximum of these two upper bounds, we can derive an upper bound on $p_{\rm acc}$ of our physically-classical verification protocol.
-We first consider the first case (i).
-Let $j^{\ast}$ such that $\tilde{m}_{0}^{(j^{\ast})}\geq m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$.
-From the Hoeffding inequality H63 , the probability of the prover not being rejected in step (d) of the $j^{\ast}$-th repetition of Protocol 2 is
-${\rm Pr}\left[\tilde{m}_{0}^{(j^{\ast})}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$ (14)
-$\leq$ ${\rm Pr}\left[m\cfrac{1+\alpha^{2}}{e^{\alpha^{2}}}-\left(m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}\right)\geq me^{-\alpha^{2}}\cfrac{\alpha^{2}}{2}\right]$
-$\leq$ ${\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),$ (15)
-where we have used $\tilde{m}_{0}^{(j^{\ast})}\geq m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$ and the fact that $m_{0}^{(j^{\ast})}+m_{1}^{(j^{\ast})}$ converges to $me^{-\alpha^{2}}(1+\alpha^{2})$ to derive the first and second inequalities, respectively FNM2 .
-Since $\tilde{m}_{0}^{(j^{\ast})}\leq me^{-\alpha^{2}}(1+\alpha^{2}/2)$ is a necessary condition to be accepted, from Eq.
-(15),
-$p_{\rm acc}\leq{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (16)
-In the second case (ii), $\tilde{m}_{0}^{(j)}<m_{0}^{(j)}+m_{1}^{(j)}$ holds, and hence the input of the I1DC protocol must include at least a single state whose actual photon number is one or zero.
-Note that we can assume that the malicious prover is not rejected in step (d) of Protocol 2 because our purpose is to derive an upper bound on $p_{\rm acc}$.
+    $m_{0}^{(j)}+m_{1}^{(j)}$ is more than the threshold value for all $1\leq j\leq N$.
+Let $p_{\rm(i)}$ and $p_{\rm(ii)}$ denote the probabilities of the first and second cases occurring, respectively.
+We also define $p_{{\rm acc}|{\rm(i)}}$ and $p_{{\rm acc}|{\rm(ii)}}$ as probabilities of the prover being accepted in the first and second cases, respectively.
+By using these notations, the acceptance probability $p_{\rm acc}$ of our physically classical verification protocol is
+$p_{\rm acc}$ $=$ $p_{{\rm acc}|{\rm(i)}}p_{\rm(i)}+p_{{\rm acc}|{\rm(ii)}}p_{\rm(ii)}\leq p_{\rm(i)}+p_{{\rm acc}|{\rm(ii)}}.$ (14)
+We derive an upper bound on $p_{\rm acc}$ by evaluating $p_{\rm(i)}$ and $p_{{\rm acc}|{\rm(ii)}}$ one by one.
+We first evaluate $p_{\rm(i)}$. From the Hoeffding inequality H63 , for all $j$,
+${\rm Pr}\left[m_{0}^{(j)}+m_{1}^{(j)}>me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$ (15)
+$=$ $1-{\rm Pr}\left[m_{0}^{(j)}+m_{1}^{(j)}\leq me^{-\alpha^{2}}\left(1+\cfrac{\alpha^{2}}{2}\right)\right]$
+$=$ $1-{\rm Pr}\left[m\cfrac{1+\alpha^{2}}{e^{\alpha^{2}}}-\left(m_{0}^{(j)}+m_{1}^{(j)}\right)\geq me^{-\alpha^{2}}\cfrac{\alpha^{2}}{2}\right]$ (16)
+$\geq$ $1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),$ (17)
+where we have used the fact that $m_{0}^{(j)}+m_{1}^{(j)}$ converges to $me^{-\alpha^{2}}(1+\alpha^{2})$ to derive the last inequality FNM2 . Therefore,
+$p_{\rm(i)}$$\leq$ $1-\left[1-{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)\right]^{N}$ (18)
+$\leq$ $N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (19)
+We next evaluate $p_{{\rm acc}|{\rm(ii)}}$.
+Let $\tilde{m}_{0}^{(j)}$ be the number of vacuum states calculated from the measurement outcomes sent by the malicious prover in step (c) of Protocol 2.
+The inequality $\tilde{m}_{0}^{(j)}<m_{0}^{(j)}+m_{1}^{(j)}$ has to hold for all $j$ to maximize $p_{{\rm acc}|{\rm(ii)}}$ because the malicious prover is definitely rejected if $\tilde{m}_{0}^{(j)}\geq m_{0}^{(j)}+m_{1}^{(j)}>me^{-\alpha^{2}}(1+\alpha^{2}/2)$.
+This implies that the input of the I1DC protocol must include at least a single state whose actual photon number is one or zero.
 We first consider the case where a single state whose actual photon number is one is included in the input of the I1DC protocol.
-Let it be the $l^{\ast}$-th input state whose polarization angle is $\sigma_{l^{\ast}}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$.
-In general, the malicious prover applies any CPTP map to the $(m-\tilde{m}_{0}^{(j)})$ input state that are used for the I1DC protocol in the case of the honest prover.
-We interpret it as that the $l^{\ast}$-th input state is converted to some quantum state by using the other $(m-\tilde{m}_{0}^{(j)}-1)$ input state as ancillary states.
+Let it be the $l^{\ast}$th input state whose polarization angle is $\sigma_{l^{\ast}}^{(j)}\in\{0,\pi/2,\pi,3\pi/2\}$.
+In general, the malicious prover applies any CPTP map to the $(m-\tilde{m}_{0}^{(j)})$ input states, which are used for the I1DC protocol in the case of an honest prover.
+We interpret it as that the $l^{\ast}$th input state is converted to some quantum state by using the other $(m-\tilde{m}_{0}^{(j)}-1)$ input states as ancillary states.
 This interpretation implies that, immediately before step (e) in Protocol 2, the prover’s state is
@@
 with some CPTP map $\mathcal{E}$.
-Since the verifier chooses each polarization angle independently in step (a) of Protocol 2, and the actual photon number of the $l^{\ast}$-th input state is one, $\mathcal{E}$ does not depend on $\sigma_{l^{\ast}}^{(j)}$ but can depend on the other input states’ polarization angles $\{\sigma_{l}^{(j)}\}_{1\leq l\leq m-\tilde{m}_{0}^{(j)},l\neq l^{\ast}}$.
+Since the verifier chooses each polarization angle independently in step (a) of Protocol 2, and the actual photon number of the $l^{\ast}$th input state is one, $\mathcal{E}$ does not depend on $\sigma_{l^{\ast}}^{(j)}$ but can depend on the other input states’ polarization angles $\{\sigma_{l}^{(j)}\}_{1\leq l\leq m-\tilde{m}_{0}^{(j)},l\neq l^{\ast}}$.
 To mimic the honest prover, in steps (e) and (f) of Protocol 2, the malicious prover generates and sends the measurement outcomes $\vec{o}^{\ (j)}\equiv\{o_{l^{\prime}}^{(j)}\}_{l^{\prime}=1}^{m-\tilde{m}_{0}^{(j)}-1}\in\{0,1\}^{m-\tilde{m}_{0}^{(j)}-1}$ by applying some additional CPTP map to the quantum state in Eq.
@@
 where $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ is the probability of outputting $\vec{o}^{\ (j)}$, and $\mathcal{E}_{\vec{o}^{\ (j)}}$ is a quantum operation to be applied when the measurement outcomes are $\vec{o}^{\ (j)}$.
-The subscript of $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ just represents that the probability, in general, depends on $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$ and does not mean that the prover’s CPTP map is constructed by using the value of $\sigma_{l^{\ast}}^{(j)}$.
-In fact, there exits a CPTP map $\mathcal{F}_{1}^{(j)}$ such that
+The subscript of $p_{\sigma_{l^{\ast}}^{(j)}}(\vec{o}^{\ (j)})$ just represents that the probability, in general, depends on $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$; it does not mean that the prover’s CPTP map is constructed by using the value of $\sigma_{l^{\ast}}^{(j)}$.
+In fact, there exists a CPTP map $\mathcal{F}_{1}^{(j)}$ such that
 $\mathcal{F}_{1}^{(j)}\left(|+_{\sigma_{l^{\ast}}^{(j)}}\rangle\langle+_{\sigma_{l^{\ast}}^{(j)}}|\right)$ (22)
@@
 for any $\sigma_{l^{\ast}}^{(j)}$.
-This is due to the fact that the verifier keeps the value of $\sigma_{l^{\ast}}^{(j)}$ private, and the actual photon number of the $l^{\ast}$-th input state is one, i.e., the information of $\sigma_{l^{\ast}}^{(j)}$ is only contained in $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$.
+This is because the verifier keeps the value of $\sigma_{l^{\ast}}^{(j)}$ private and the actual photon number of the $l^{\ast}$th input state is one, i.e., the information of $\sigma_{l^{\ast}}^{(j)}$ is only contained in $|+_{\sigma_{l^{\ast}}^{(j)}}\rangle$.
 For any $\vec{o}^{\ (j)}$ and $\{\sigma_{l}^{(j)}\}_{l\neq l^{\ast}}$, there exist $\sigma\in\{0,\pi/2,\pi,3\pi/2\}$ and $c\in\{1,-1\}$ such that
@@
 We next consider the case where a vacuum state is included in the input of the I1DC protocol.
-Let it be originated from the $k^{\ast}$-th coherent state whose polarization angle is $\theta_{k^{\ast}}^{(j)}$.
+Let the vacuum state be originated from the $k^{\ast}$th coherent state whose polarization angle is $\theta_{k^{\ast}}^{(j)}$.
 Since the verifier selects each polarization angle independently, the malicious prover’s state $\rho^{(j)}$ at the end of Protocol 2 does not depend on $\theta_{k^{\ast}}^{(j)}$ and hence $\varphi^{(j)}$.
@@
 This quantum state can also be prepared at the end of step 1 in Protocol 1.
-By combining the above arguments, when the deviation of the prover is independent in each repetition, the malicious prover’s state after the $N$-th repetition in Protocol 2 is
+By combining the above arguments, when the deviation of the prover is independent in each repetition, the malicious prover’s state after the $N$th repetition in Protocol 2 is
 $\cfrac{1}{4^{N}}\sum_{\{\varphi^{(j)}\}_{j=1}^{N}}\bigotimes_{j=1}^{N}\mathcal{F}^{(j)}\left(|\psi_{V}^{(j)}\rangle\langle\psi_{V}^{(j)}|\right)$ (28)
@@
 where $\mathcal{F}^{(j)}\in\{\mathcal{F}_{1}^{(j)},\mathcal{F}_{2}^{(j)}\}$ for all $1\leq j\leq N$.
-As shown in Appendix B, the similar argument holds even when the prover’s attack is collective, i.e., the prover simultaneously handles all photons in all repetitions to deceive the verifier.
-Therefore, $p_{\rm acc}\leq 1-b/2$.
-From this upper bound and Eq.
-(16), when the prover is malicious,
-$p_{\rm acc}\leq{\rm max}\left\{{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),1-\cfrac{b}{2}\right\}.$ (27)
+As shown in Appendix B, a similar argument holds even when the prover’s attack is collective, i.e., the prover simultaneously handles all photons in all repetitions to deceive the verifier.
+Therefore, $p_{{\rm acc}|{\rm(ii)}}\leq 1-b/2$.
+From this upper bound and Eqs.
+(14) and (19), when the prover is malicious,
+$p_{\rm acc}\leq 1-\cfrac{b}{2}+N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right).$ (30)
 […]
 $1-\cfrac{a}{2}-N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)$ (31)
-$-{\rm max}\left\{{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right),1-\cfrac{b}{2}\right\}$
-$\geq$ $1-\cfrac{a}{2}-\cfrac{1}{2Nf(|x|)}-\left(1-\cfrac{b}{2}\right)$
+$-\left[1-\cfrac{b}{2}+N{\rm exp}\left(-\cfrac{me^{-2\alpha^{2}}\alpha^{4}}{2}\right)\right]$
+$\geq$ $\cfrac{b-a}{2}-\cfrac{1}{2Nf(|x|)}$
 $\geq$ $\cfrac{N-1}{2Nf(|x|)},$ (32)
-which is the inverse of a polynomial in $|x|$. Thus, our physically-classical verification protocol efficiently distinguishes the honest and malicious provers.
+which is the inverse of a polynomial in $|x|$. Thus, our physically classical verification protocol efficiently distinguishes the honest and malicious provers.
 IV Conclusion & discussion
-We have proposed an efficient verification protocol that removes the necessity of the communication of qubits.
+We have proposed an efficient verification protocol that removes the necessity of qubit communication.
 Since all apparatuses required for the verifier are a telecom-band laser with linear optical elements and a classical computer, our results would facilitate the realization of verifiable quantum computers.
 To further improve the practicality of our protocol, we discuss the phase randomization implemented for the coherent state.
-In several papers (e.g., Ref.
-CZLM15 ), the difficulty of the continuous phase randomization was pointed out.
+Several papers (e.g., Ref.
+CZLM15 ) have pointed out the difficulty of the continuous phase randomization.
 By calculating the fidelity between the continuous-phase-randomized (see Eq.
@@
 $\geq$ $1-2mN\left(\cfrac{e}{R-1}\right)^{R-1}\equiv F_{\rm min}.$ (39)
-Let $p_{\rm acc}$ and $q_{\rm acc}$ be the acceptance probabilities of our physically-classical verification protocol with continuous-phase-randomized and discrete-phase-randomized coherent states, respectively.
+Let $p_{\rm acc}$ and $q_{\rm acc}$ be the acceptance probabilities of our physically classical verification protocol with continuous-phase-randomized and discrete-phase-randomized coherent states, respectively.
 Eq.
-(29) implies that $|p_{\rm acc}-q_{\rm acc}|\leq(N-1)/[4Nf(|x|)]$ is sufficient to that our protocol correctly distinguishes the honest and malicious provers.
+(32) implies that $|p_{\rm acc}-q_{\rm acc}|\leq(N-1)/[4Nf(|x|)]$ is sufficient for our protocol to correctly distinguish between honest and malicious provers.
 From Ref.
 HQ12 that studies the fidelity and trace distance for any quantum states in an infinite-dimensional separable complex Hilbert space, $|p_{\rm acc}-q_{\rm acc}|\leq\sqrt{1-F_{\rm min}}$, and hence using $R\geq e^{2}+1$ results in $R=\lceil\log{[32mN^{3}f(|x|)^{2}/(N-1)^{2}]}+1\rceil$.
-Since $m=\lceil 2e^{2}\log{(2N^{2}f(|x|))}\rceil$, this calculation shows that a logarithmic of logarithmic number of classical bits is sufficient for the phase randomization.
+Since $m=\lceil 2e^{2}\log{(4N^{2}f(|x|))}\rceil$, this calculation shows that a log of log number of classical bits is sufficient for the phase randomization.
 The RBSP protocol DKL12 was improved or modified in Refs.
@@
 As another direction to improve our protocol, it would be interesting to consider their applicability to the verification of quantum computation.
-Furthermore, as with the implementation security ZNC24 of QKD, it would be important to devise the verification protocols under several imperfections such as the channel loss and noises and the correlation between coherent states.
+Furthermore, as with the implementation security ZNC24 of QKD, it would be important to devise the verification protocols under several imperfections such as channel loss and noises and the correlation between coherent states.
 An efficient way KLMO24 that removes trusted quantum state preparations and measurements from verification protocols may be useful for this purpose.
-In this paper, we propose a physically-classical verification protocol by combining the protocols in Refs.
+In this paper, we have proposed a physically classical verification protocol by combining the protocols in Refs.
 M20 and DKL12 .
@@
 Although Ref.
-M18 seems to implicitly assume the perfect blindness (i.e., the perfect security), the blind quantum computing protocol with coherent states proposed in Ref.
-DKL12 does not satisfy the perfect blindness.
-This is why it would not be directly applied to Ref.
-DKL12 to obtain a physically-classical verification protocol.
+M18 seems to implicitly assume perfect blindness (i.e., perfect security), the blind quantum computing protocol with coherent states proposed in Ref.
+DKL12 does not have perfect blindness.
+This is why it could not be directly applied to Ref.
+DKL12 to obtain a physically classical verification protocol.
 When we apply Ref.
@@
 1 with a constant number of rounds can be constructed, then ${\sf BQP}$ is contained in the third level of the polynomial hierarchy FHM18 .
-This set containment is considered to be unlikely from an oracle separation between ${\sf BQP}$ and ${\sf PH}$ RT22 .
-It would be interesting to devise physically-classical verification protocols by modifying other existing verification protocols (e.g., Ref. B18 ) and then combining it with the RBSP protocol DKL12 .
+This set containment is considered to be unlikely due to an oracle separation between ${\sf BQP}$ and ${\sf PH}$ RT22 .
+It would be interesting to devise physically classical verification protocols by modifying other existing verification protocols (e.g., Ref.
+B18 ) and then combining them with the RBSP protocol DKL12 .
 ACKNOWLEDGMENTS
@@
 Appendix B: Malicious prover’s final state in Protocol 2 for collective attacks
-In this appendix, we derive the malicious prover’s quantum state after the $N$-th repetition in Protocol 2 under any collective attack.
-Since we consider the second case (ii) where the number $\tilde{m}_{0}^{(j)}$ of vacuum states is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$, the I1DC protocol in the $j$-th repetition has a state whose actual photon number is zero or one as an input.
-Let it be the $l^{\ast}_{j}$-th input state whose polarization angle is $\sigma_{l^{\ast}_{j}}^{(j)}$.
+In this appendix, we derive the malicious prover’s quantum state after the $N$th repetition in Protocol 2 under any collective attack.
+Since we consider case (ii), we can assume that the number $\tilde{m}_{0}^{(j)}$ of vacuum states is less than $m_{0}^{(j)}+m_{1}^{(j)}$ for all $1\leq j\leq N$.
+Therefore, the I1DC protocol in the $j$th repetition has a state whose actual photon number is zero or one as an input.
+Let it be the $l^{\ast}_{j}$th input state whose polarization angle is $\sigma_{l^{\ast}_{j}}^{(j)}$.
 For simplicity, we define $\vec{\sigma}\in\{0,\pi/2,\pi,3\pi/2\}^{(m-1)N}$ as the string of all the polarization angles except for $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$.
@@
 where $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ is the probability of outputting $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$, and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ is a quantum operation to be applied when the measurement outcomes are $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$.
-The subscript $\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ just represents that the probability can depend on each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ if the $l^{\ast}_{j}$-th input qubit of the I1DC protocol in the $j$-th repetition is not a vacuum state, and it does not mean that the prover’s CPTP map is constructed by using the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$.
-Another subscript $\vec{\sigma}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ means that they may depend on $\vec{\sigma}$ because the corresponding coherent states may include more than one photon.
-Since the verifier keeps the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$ private, and the actual photon number of each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is one or zero, as with Eq.
+The subscript $\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)}$ of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ just represents that the probability can depend on each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ if the $l^{\ast}_{j}$th input qubit of the I1DC protocol in the $j$th repetition is not a vacuum state; it does not mean that the prover’s CPTP map is constructed by using the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$.
+The other subscript, $\vec{\sigma}$, of $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ and $\mathcal{E}_{\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)},\vec{\sigma}}$ means that they may depend on $\vec{\sigma}$ because the corresponding coherent states may include more than one photon.
+Since the verifier keeps the values of $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$ private and the actual photon number of each $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is one or zero, as with Eq.
 (22), there exists a CPTP map $\mathcal{F}$ such that
@@
 for any $\{\sigma_{l^{\ast}_{j}}^{(j)}\}_{j=1}^{N}$.
-Remember that when the actual photon number of $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is zero, the probability $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ does not depend on $\sigma_{l^{\ast}_{j}}^{(j)}$.
+Recall that when the actual photon number of $|+_{\sigma_{l^{\ast}_{j}}^{(j)}}\rangle$ is zero, the probability $p_{\sigma_{l^{\ast}_{1}}^{(1)},\sigma_{l^{\ast}_{2}}^{(2)},\ldots,\sigma_{l^{\ast}_{N}}^{(N)},\vec{\sigma}}(\vec{o}^{\ (1)},\vec{o}^{\ (2)},\ldots,\vec{o}^{\ (N)})$ does not depend on $\sigma_{l^{\ast}_{j}}^{(j)}$.
 For any $\vec{\sigma}$ and $\{\vec{o}^{\ (j)}\}_{j=1}^{N}$, there exist $\{\theta_{j}\}_{j=1}^{N}\in\{0,\pi/2,\pi,3\pi/2\}^{N}$ and $\{c_{j}\}_{j=1}^{N}\in\{1,-1\}^{N}$ such that
@@
 $=$ $\cfrac{1}{4^{N}}\sum_{\varphi^{(1)},\varphi^{(2)},\ldots,\varphi^{(N)}}\mathcal{F}\left(|\psi_{V}\rangle\langle\psi_{V}|\right).$ (59)
-Since this quantum state can be prepared at the end of step 1 in Protocol 1, the inequality $p_{\rm acc}\leq 1-b/2$ holds.
+Since this quantum state can be prepared at the end of step 1 in Protocol 1, the inequality $p_{{\rm acc}|{\rm(ii)}}\leq 1-b/2$ holds.